Suppose among 32 people who save paper or bottles (or both) for recycling,


there are 30 who save paper and 14 who save bottles. Find the number m of


people who:


(a) save both; (b) save only paper; (c) save only bottles.

1
Expert's answer
2021-12-20T16:22:54-0500
"N(P\cup B)=N(P)+N(B)-N(P\cap B)"

(a)

"N(P\cap B)=30+14-32=12"

(b)

"N(only\ P)=N(P)-N(P\cap B)=30-12=18"

(c)

"N(only\ B)=N(B)-N(P\cap B)=14-12=2"